OpenStudy (kobeni-chan):

Connor went to New York for vacation. He spent 3 nights at a hotel and rented a car for 4 days. Jillian stayed at the same hotel, but spent 4 nights and rented a car for 5 days from the same company. If Connor paid $675 and Jillian paid $875, how much did one night at the hotel cost?

OpenStudy (kobeni-chan):

I set up equations as 3h + 4c = $675, and 4h + 5c = $875, but I don't rememebr what to do from there.

OpenStudy (kobeni-chan):

*remember

OpenStudy (animal_lover36):

Connor's = $675 = H x 3 nights + R x 4 days Jillian's = $875 = H x 4 nights + R x 5 days ($675 = H x 3 nights + R x 4 days) x (5) = $3,375 = 15 nights x H + 20 days x R ($875 = H x 4 nights + R x 5 days) x (-4) = - $3,500 = -16 nights x H + -20 days x R $3,375 = 15 nights x H + 20 days x R - $3,500 = -16 nights x H + -20 days x R H = $125

OpenStudy (kobeni-chan):

Oh ok :) thank you!

OpenStudy (animal_lover36):

np :)